### v1

# QuiZSF: An efficient data-model interaction framework for zero-shot time-series forecasting

Abstract
Time series forecasting has become increasingly important to empower diverse applications with streaming data. Zero-shot time-series forecasting (ZSF), particularly valuable in data-scarce scenarios, such as domain transfer or forecasting under extreme conditions, is difficult for traditional models to deal with. While time series pre-trained models (TSPMs) have demonstrated strong performance in ZSF, they often lack mechanisms to dynamically incorporate external knowledge. Fortunately, emerging retrieval-augmented generation (RAG) offers a promising path for injecting such knowledge on demand, yet they are rarely integrated with TSPMs. To leverage the strengths of both worlds, we introduce RAG into TSPMs to enhance zero-shot time series forecasting. In this paper, we propose QuiZSF (Quick Zero-Shot Time Series Forecaster), a lightweight and modular framework that couples efficient retrieval with representation learning and model adaptation for ZSF. Specifically, we construct a hierarchical tree-structured ChronoRAG Base (CRB) for scalable time-series storage and domain-aware retrieval, introduce a Multi-grained Series Interaction Learner (MSIL) to extract fine- and coarse-grained relational features, and develop a dual-branch Model Cooperation Coherer (MCC) that aligns retrieved knowledge with two kinds of TSPMs: Non-LLM based and LLM based. Compared with contemporary baselines, QuiZSF, with Non-LLM based and LLM based TSPMs as base model, respectively, ranks Top1 in 75% and 87.5% of prediction settings, while maintaining high efficiency in memory and inference time.
1 Introduction
Time series forecasting (TSF) is crucial for data-driven applications in dynamic systems, ranging from weather prediction, electricity consumption estimation to economic dynamic forecasting. Despite the excellent performance of many TSF studies, they face challenges in data-scarce situations wu2022timesnet ; gruver2024large ; jin2023time , such as extreme weather prediction and cross-domain adaptation, where limited historical data hinder accurate forecasting. Since it is impossible to summarize all series patterns within a system liutimer , zero-shot time forecasting (ZSF), which exploits available patterns to infer the trends and future evolution of new time series, has become a prevailing and essential research topic ekambaram2024tiny .
Recently, there has been growing interest in developing general times series pre-trained models (TSPMs) for ZSF. These models aim to transfer learned knowledge to previously unseen target time series (TS) datasets, similar to the success achieved in natural language processing (NLP) and computer vision (CV). TSPMs can be mainly classified into two categories: Non-LLM based TSPMs and LLM based TSPMs (ekambaram2024tiny ). The former pre-trains models on multiple TS datasets, represented by models like Moment (goswami2024moment ), TTM (ekambaram2024tiny ), TimesFM (das2024decoder ), Moirai (woo2024unified ), and Lag-LLaMA (rasul2023lag ); the latter uses pre-trained large language models (LLMs) for TSF, treating the prediction task as a form of cross-domain transfer learning, represented by models such as LLMTime (gruver2024large ), Time-LLM (jin2023time ), and GPT4TS (zhou2023one ). Both types of TSPMs have demonstrated good performance in ZSF.
However, using TSPMs for ZSF faces two main issues: (1) Real-world data is highly dynamic, but TSPMs cannot update their knowledge in real-time. Fine-tuning on new data is computationally expensive and disrupts the plug-and-play goal of ZSF tan2024language . The issue is how to enable models to learn dynamically updated knowledge without incurring excessive costs while maintaining a plug-and-play approach. (2) Time series often exhibit structural similarity tire2024retrieval , which is especially valuable in sparse-data ZSF settings. The issue is how to leverage inter- and intra-domain data with patterns similar to the target sequence to enhance prediction accuracy.
To address these issues, we propose integrating TSPMs with retrieval-augmented generation (RAG). For issue (1), RAG provides efficient access to updated external knowledge, avoiding costly fine-tuning and preserving the plug-and-play nature of ZSF. For issue (2), RAG retrieves relevant sequences from inter- and intra-domain datasets, offering contextual information such as periodic trends, spikes, or transitions. This enhances model generalization and prediction accuracy in sparse-data scenarios. As shown in Figure 1, RAG identifies similar TS in other datasets to improve ZSF for ETTm1.
Figure 1: Motivation. Time series across domains often exhibit similar temporal patterns, which can be retrieved and reused as auxiliary knowledge.
Nevertheless, integrating external retrieval with TSFMs introduces three fundamental challenges: (i) Efficient and Domain-Sensitive Storage and Retrieval: How to structure and index millions of TS for fast, relevant, and cross-domain retrieval remains a foundational bottleneck in large-scale systems. (ii) Multi-Level Feature Extraction: Retrieved sequences vary in scale, domain, and noise level. Designing a lightweight yet expressive architecture to extract multi-level interaction patterns between target and retrieved series is essential for generalization. (iii) Modality-Aligned Representation Integration: TSPMs are categorized into Non-LLM based and LLM based types. The former requires seamless fusion of numerical features, while the latter demands coherent transformation of retrieved data into language prompts. A unified adapter must bridge these modalities while supporting gradient-based optimization.
To address these challenges, we propose the Quick Zero-Shot Time Series Forecaster (QuiZSF), which, to the best of our knowledge, is the first efficient data–model interaction framework for ZSF. QuiZSF integrates scalable storage and retrieval, informative sequence representation learning, and modality-aware model adaptation. Our main contributions are summarized as follows:
  • We construct the ChronoRAG Base (CRB), a hierarchical tree-structured temporal database, and propose a Hybrid and Hierarchical TS Retrieval (HHTR) strategy that enables fast, scalable, and domain-sensitive retrieval across large-scale time-series databases.
  • We propose the Multi-grained Series Interaction Learner (MSIL), a lightweight yet expressive module that extracts both fine-grained dependencies and coarse-grained patterns from retrieved sequences to enhance target sequence understanding.
  • We present the Model Cooperation Coherer (MCC), a dual-branch adapter that integrates retrieved features into both Non-LLM and LLM-based TSPMs. MCC constructs residual-semantic fusion inputs for numerical models and structured prompts for language models, while supporting end-to-end optimization across modalities.
  • We conduct experiments using Non-LLM and LLM based TSPMs as base models, separately. The former achieves first place in 75% of ZSF settings, while the latter ranks first in 87.5% of the settings, all while maintaining high efficiency in memory usage and inference speed.
2 Related Work
2.1 TSPMs for Zero-shot Time Series Forecasting
Recent advancements in TSPMs for ZSF have garnered significant attention. These models can be broadly categorized into two types. The first is pre-trained models designed specifically for TSF. These models learn generalizable temporal representations from large-scale TS datasets.Representative methods include Moment goswami2024moment , TTM ekambaram2024tiny , TimesFM das2024decoder , Moirai woo2024unified , and Lag-LLaMA rasul2023lag . For example, TTM adopts a lightweight mixer-style architecture and demonstrates strong zero-shot performance across multiple domains. The second type leverages pre-trained large language models (LLMs) by framing time-series forecasting as a form of cross-domain transfer learning. These models, including LLMTime gruver2024large , Time-LLM jin2023time , and GPT4TS zhou2023one , transform numerical sequences into natural language prompts, enabling ZSF.
Despite promising results, these models rely heavily on large-scale pre-training or fine-tuning across diverse datasets. They often lack scalability and the ability to incorporate real-time, open-world knowledge. As illustrated in Figure 1, time series from similar domains frequently share common temporal patterns, which can serve as valuable references in low-resource settings. With the advancement of retrieval techniques, integrating external sequence-level knowledge into pre-trained models enables dynamic forecasting, mitigates hallucinations, and constrains predictions within realistic boundaries. Enhancing TSPMs with retrieval capabilities has therefore emerged as a promising direction for improving zero-shot forecasting.
2.2 Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) combines models with information retrieval to enhance model performance by leveraging external knowledge. Early work, such as REALM guu2020retrieval , introduces retrieval-augmented models by incorporating unsupervised retrieval modules, allowing models to access relevant passages during generation. RAG lewis2020retrieval further advances this by combining retrieval and generation in an end-to-end framework, improving tasks like open-domain question answering. Subsequent approaches, including T5+RAG lewis2020retrieval and Fusion-in-Decoder (FiD) izacard2020leveraging , enhance the fusion of retrieved information, achieving better coherence and relevance. Models like RETRO borgeaud2022improving have pushed the boundaries by incorporating retrieval at both pre- and post-generation stages, further improving output quality. Applying RAG to time series forecasting offers promising benefits. Recent studies jing2022retrieval ; yeh2023efficient ; tire2024retrieval have introduced retrieval mechanisms into forecasting. However, they often overlook the cost and scalability of external databases in large-scale settings, leading to inefficiencies. While TimeRAF zhang2024timeraf explores zero-shot forecasting with retrieval, it lacks a clear mechanism for modeling interactions between pre-trained models and retrieved sequences. These pioneering studies highlight the potential value of combining pre-trained models with retrieval processes.
3 Preliminaries and Problem Definition
We focus on zero-shot forecasting (ZSF) with Time Series Pre-trained Models (TSPMs), which can be categorized into Non-LLM based (processing numerical inputs) and LLM based (operating on language inputs). We conduct separate ZSF formulations to accommodate their distinct modalities.
3.1 Retrieval-Augmented Non-LLM based TSPMs
Let $\mathcal{G}=\{\mathcal{G}_{1},\dots,\mathcal{G}_{K}\}$ denote multiple distinct TS domains, where each domain $\mathcal{G}_{k}$ contains series $\mathcal{G}_{k}=\{\mathbf{X}^{k}\mid x^{k}_{1},\dots,x^{k}_{T}\}$. These series are compiled into a dynamic auxiliary database $\mathcal{D}$.
Given a target sequence $\mathbf{X}^{T}$, we retrieve the top-$K$ relevant auxiliary series $\mathbb{X}^{R}=\{\mathbf{X}^{R}_{1},\dots,\mathcal{X}^{R}_{K}\}\subset\mathcal{D}$. Since Non-LLM models do not support textual prompts, the target and retrieved series are fused into a new representation, $\widehat{\mathbf{X}}^{T}=\mathcal{F}_{\text{N}}^{*}(\mathbf{X}^{T},\mathbb{X}^{R}),\quad\widehat{\mathbf{Y}}^{T}=\mathcal{M}_{1}^{*}(\widehat{\mathbf{X}}^{T})$, where $\mathcal{M}_{1}$ is the modified learning framework with Non-LLM based TSPM.
3.2 Retrieval-Augmented LLM based TSPMs
In contrast, for LLM based TSPMs, we first process the target sequence and retrieved auxiliary series to produce both a structured input and a corresponding textual prompt:
$\widehat{\mathbf{X}}^{T},\mathcal{P}=\mathcal{F}_{\text{L}}^{*}(\mathbf{X}^{T},\mathbb{X}^{R}),\quad\widehat{\mathbf{Y}}^{T}=\mathcal{M}_{2}^{*}(\widehat{\mathbf{X}}^{T}\mid\mathcal{P}),$ (1)
where $\mathcal{M}_{2}$ is the modified learning framework with LLM based TSPM and $\mathcal{P}$ is the generated prompt fusing the retrieved knowledge.
Figure 2: Overview of QuiZSF.
4 Methodology
4.1 Framework overview
Our zero-shot learning framework consists of three major components, as illustrated in Figure 2. 1) a ChronoRAG Base with efficient hierarchical retrieval, 2) a Multi-grained Series Interaction Learner captures series-level interactions between the target and retrieved series, offering valuable information for zero-shot forecasting, 3) a Model Cooperation Coherer enables modality-aligned integration between learned representations and both Non-LLM based and LLM based TSPMs.
4.2 ChronoRAG Base
ChronoRAG Base (CRB) is constructed from 27 time series datasets across seven domains: Energy, Health, IoT, Nature, Web, Transport, and Environment. It integrates five public data sources: Monash godahewa2021monash , UCR dau2019ucr , TSER tan2021time , TDBrain van2022two , and UTSD liutimer . These datasets span diverse sampling granularities, ranging from daily to minute-level resolutions. We provide three CRB versions: Small, Medium, and Large, containing 34M, 48M, and 143M time points, respectively. All versions cover the same domains, and smaller versions are strict subsets of the larger ones. Detailed statistics of CRB-Large are shown in Table 5. Two core components support CRB as follows.
Data protocol for ChronoRAG. To ensure consistency and scalability, we design a unified data protocol including sliding-window segmentation, linear interpolation for missing values, channel-independent processing for dimensional alignment, and standardized metadata. All sequences are stored in the ARROW format peltenburg2019fletcher for efficient access. Further Details are provided in Appendix A.1.
Hierarchical Series Tree. We implement a hierarchical tree structure to support fast indexing and retrieval. The base is partitioned by domain and clustered using $k$-means to enable efficient approximate search and incremental updates. Further Details are provided in Appendix 6.
4.3 Hybrid and Hierarchical Time-series Retrieval
To support both high-accuracy and scalable search, we propose a hybrid and hierarchical TS retrieval (HHTR) strategy. It combines local domain-specific matching and global prototype comparison, leveraging the hierarchical index structure built in CRB. The strategy is shown in Figure 3 (a).
Top-K Series Retrieval. Given a query time series $\mathbf{X}_{T}$, the objective is to retrieve its most relevant neighbors from the prototype database. This is a classical nearest-neighbor retrieval task, tailored to the time-series domain through hybrid matching strategies. If the domain of $\mathbf{X}_{T}$ is known and exists in CRB, we apply a combination of local and global prototype matching. The final Top-$K$ set is computed as,
$\text{Top}\;\;K=\rho\cdot\text{Top}\;\;K_{\text{local}}+(1-\rho)\cdot\text{Top}\;\;K_{\text{global}},$ (2)
where $\rho\in[0,1]$ controls the balance between local and global contributions. If the domain is not identified or not present in CRB, retrieval is performed across all stored cluster prototypes,
$\text{Top}\;\;K=\text{Top}\;\;K_{\text{global}},$ (3)
This hybrid design enhances both retrieval accuracy and generalization, especially in settings with noisy or ambiguous domain information.
Distance Metric Design. To quantify similarity between the query $\mathbf{X}_{T}$ and a candidate sequence $\mathbf{X}_{i}$, we define a compound similarity score:
$\text{Sim}(\mathbf{X}_{T},\mathbf{X}_{i})=\cos(\mathbf{X}_{T},\mathbf{X}_{i})+\frac{1}{\text{dist}(\mathbf{X}_{T},\mathbf{X}_{i})},$ (4)
where $\cos(\cdot,\cdot)$ denotes cosine similarity and $\text{dist}(\cdot,\cdot)$ is the Euclidean distance. This dual metric emphasizes trend alignment through cosine similarity while capturing geometric proximity via Euclidean distance.
Figure 3: (a) HHTR: Domain-aware and global retrieval via a hierarchical index. (b) MSIL: Interaction and average patterns extracted from retrieved sequences are fused with the target via cross-attention.
4.4 Multi-grained Series Interaction Learner
The core challenge after retrieving relevant series lies in effectively integrating them with the target series to enhance forecasting. Simply concatenating or averaging the retrieved series can lead to suboptimal results due to heterogeneity across domains. To this end, we propose the Multi-grained Series Interaction Learner (MSIL), which is designed to extract robust and informative representations by modeling fine-grained interactions and global trends simultaneously.
MSIL is motivated by three considerations: (1) retrieved series may come from different domains with distinct dynamics; (2) fine-grained dependencies across series are often crucial for prediction; and (3) domain-specific context and global knowledge should be fused in a unified representation. As illustrated in Figure 3 (b), MSIL achieves this by computing two representative patterns, namely an interaction pattern and an average pattern, which are then fused with the target sequence through a cross-attention mechanism.
Given a target time series $\mathbf{T}\in\mathbb{R}^{N\times D}$ with $N$ time steps and $D$ channels, and a set of retrieved series $\{\mathbf{S}_{1},\dots,\mathbf{S}_{n}\}$, we first divide the retrieved set into same-domain and cross-domain subsets:
$\mathbb{S}_{\text{same}}=\{\mathbf{S}_{i}\mid\text{domain}(\mathbf{S}_{i})=\text{domain}(\mathbf{T})\},\quad\mathbb{S}_{\text{cross}}=\{\mathbf{S}_{j}\mid\text{domain}(\mathbf{S}_{j})\neq\text{domain}(\mathbf{T})\}.$ (5)
To ensure numerical consistency, we normalize both target and retrieved sequences using the scaler module associated with their respective base models,
$\mathbf{T}^{\text{norm}},\text{loc}_{T},\text{scale}_{T}$ $=\text{scaler}(\mathbf{T}),$ $\mathbf{S}_{i}^{\text{norm}},\text{loc}_{S_{i}},\text{scale}_{S_{i}}$ $=\text{scaler}(\mathbf{S}_{i}),\quad\forall i,$ (6)
with
$\mathbf{T}^{\text{norm}}=\frac{\mathbf{T}-\text{loc}_{T}}{\text{scale}_{T}},\quad\mathbf{S}_{i}^{\text{norm}}=\frac{\mathbf{S}_{i}-\text{loc}_{S_{i}}}{\text{scale}_{S_{i}}}.$ (7)
Based on the normalized retrieved sequences, MSIL computes two complementary patterns: 1) Interaction Pattern ($\mathbf{P}_{\text{int}}$) captures fine-grained dependencies via element-wise product, followed by a non-linear projection. 2) Average Pattern ($\mathbf{P}_{\text{avg}}$) encodes global trends through mean pooling and transformation. These are defined as,
$\mathbf{P}_{\text{int}}$ $=\text{MLP}_{1}\left(\frac{\prod_{i=1}^{n}\mathbf{S}_{i}^{\text{norm}}}{\left\|\prod_{i=1}^{n}\mathbf{S}_{i}^{\text{norm}}\right\|}\right),$ $\quad\mathbf{P}_{\text{avg}}$ $=\text{MLP}_{2}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{S}_{i}^{\text{norm}}\right).$ (8)
To fuse these patterns with the target sequence, we use a multi-grained cross-attention mechanism, where the target serves as query and the patterns as key/value,
$\mathbf{Q}=\mathbf{W}_{q}\mathbf{T}^{\text{norm}},\quad\mathbf{K}=\mathbf{W}_{k}\mathbf{P}_{\text{avg}},\quad\mathbf{V}=\mathbf{W}_{v}\mathbf{P}_{\text{int}},$ (9)
$\mathbf{R}_{\text{fused}}=\text{Softmax}\left(\frac{\mathbf{Q}\cdot\mathbf{K}^{T}}{\sqrt{d}}\right)\cdot\mathbf{V}.$ (10)
The resulting representation $\mathbf{R}_{\text{fused}}$ combines domain-aware alignment and trend-aware feature fusion. MSIL enables rich interaction modeling between the target and retrieved series, and its multi-granular design improves generalization, especially in zero-shot forecasting scenarios.
4.5 Model Cooperation Coherer
In retrieval-augmented zero-shot forecasting, it is crucial to effectively connect retrieved knowledge with diverse TSPMs. These models fall into two categories: Non-LLM-based (numerical input) and LLM-based (textual input), each requiring tailored integration strategies. To fully leverage the representations produced by MSIL, we design a unified cooperation mechanism with two branches: one for numerical models and one for language models, both supporting feedback-driven optimization.
Numerical Coherer for Non-LLM based TSPMs. For numerical TS pre-trained models, we apply a residual connection to fuse the normalized target sequence $\mathbf{T}^{\text{norm}}$ with the MSIL-fused representation $\mathbf{R}_{\text{fused}}$, then feed it into the forecasting model,
$\widehat{\mathbf{T}}=\mathcal{F}_{\text{N}}(\mathbf{R}_{\text{fused}},\mathbf{T}^{\text{norm}}),$ (11)
where $\mathcal{F}_{\text{N}}$ is a residual module he2016deep that enhances expressiveness and mitigates gradient vanishing. The output $\widehat{\mathbf{T}}$ is the forecasted sequence.
Language Coherer for LLM based TSPMs. Language models operate exclusively on text, making direct use of numeric features infeasible. To bridge this modality gap, we convert MSIL outputs ($\mathbf{P}_{\text{int}},\mathbf{P}_{\text{avg}},\mathbf{T}^{\text{norm}}$) into structured textual summaries. These summaries, along with an instruction-style prompt, guide the language model in generating forecasting outputs. See Appendix B and Figure 8 for prompt construction details.
MMD-based Optimization Strategy. To further align the predicted and ground truth distributions, we introduce a regularization term based on Maximum Mean Discrepancy (MMD) borgwardt2006integrating . Formally, given predicted sequences $\widehat{\mathbf{T}}\sim\mathcal{P}_{\widehat{T}}$ and ground truth sequences $\mathbf{T}_{\text{true}}\sim\mathcal{P}_{T}$, the MMD is defined as:
$\text{MMD}^{2}(\mathcal{P}_{\widehat{T}},\mathcal{P}_{T})=\left\|\mathbb{E}_{\widehat{\mathbf{T}}}[\phi(\widehat{\mathbf{T}})]-\mathbb{E}_{\mathbf{T}_{\text{true}}}[\phi(\mathbf{T}_{\text{true}})]\right\|_{\mathcal{H}}^{2},$ (12)
where $\phi(\cdot)$ is a feature map that embeds time series into a Reproducing Kernel Hilbert Space (RKHS) $\mathcal{H}$, and $\|\cdot\|_{\mathcal{H}}$ denotes the RKHS norm. Using the kernel trick with $k(\mathbf{x},\mathbf{y})=\langle\phi(\mathbf{x}),\phi(\mathbf{y})\rangle$, MMD can be computed as:
$\text{MMD}^{2}=\mathbb{E}_{\widehat{\mathbf{T}},\widehat{\mathbf{T}}^{\prime}}[k(\widehat{\mathbf{T}},\widehat{\mathbf{T}}^{\prime})]+\mathbb{E}_{\mathbf{T}_{\text{true}},\mathbf{T}_{\text{true}}^{\prime}}[k(\mathbf{T}_{\text{true}},\mathbf{T}_{\text{true}}^{\prime})]-2\mathbb{E}_{\widehat{\mathbf{T}},\mathbf{T}_{\text{true}}}[k(\widehat{\mathbf{T}},\mathbf{T}_{\text{true}})].$ (13)
This formulation encourages the distribution of predicted sequences to match that of the true sequences. We integrate this into the training objective as a regularizer:
$\mathcal{L}=\mathcal{L}_{\text{MSE}}(\widehat{\mathbf{T}},\mathbf{T}_{\text{true}})+\lambda\cdot\text{MMD}^{2}(\widehat{\mathbf{T}},\mathbf{T}_{\text{true}}),$ (14)
where $\lambda$ is a balancing hyperparameter. This loss is backpropagated to update the learnable parameters, enhancing distributional alignment between prediction and truth.
5 Experiments
5.1 Experiments Setups
For evaluation datasets, we use five public datasets: ETTh1, ETTh2, ETTm1, ETTm2, and Weather, which are widely used in prior state-of-the-art works nie2022time ; jin2023time ; zhou2023one . Standard error metric is MSE. For CRB selection, our ChronoRAG Base comes in three versions: CRB-Small, CRB-Medium, and CRB-Large, with details in Appendix A.3. Given the balance between computing time and experimental results, we select CRB-Medium as the RAG Base. For model comparison, we evaluate against thirteen state-of-the-art forecasting methods, which can be classified into the following categories: (a) Non-LLM based TSPMs: TTM ekambaram2024tiny , Moirai woo2024unified , and TimesFM das2024decoder ; (b) LLM based TSPMs: TimeLLM jin2023time , LLMTime gruver2024large , and GPT4TS zhou2023one ; (c) Other architectures: iTransformer liu2023itransformer , Crossformer zhang2023crossformer , DLinear zeng2023transformers , TimesNet wu2022timesnet , PatchTST nie2022time , TiDE das2023long , and FEDformer zhou2022fedformer .
5.2 Implementation Details
In the time series domain, zero-shot forecasting refers to evaluating models on unseen datasets without direct supervision or fine-tuning. Given the two types of TSPMs, current ZSF setups are categorized accordingly: the multi-source generalization setup for Non-LLM based models, and the single-source transfer setup for LLM based models. To ensure comprehensive evaluation, we adopt both setups in our experiments, resulting in QuiZSFT and QuiZSFL, respectively.
Multi-source generalization zero-shot setup (for Non-LLM based TSPMs) This setup trains on diverse source datasets and enables zero-shot forecasting by directly applying the model to unseen target datasets. This approach is widely used in Non-LLM based TSPMs like TTM ekambaram2024tiny , TimesFM das2024decoder , and Moirai woo2024unified . We use this setup for QuiZSFT, with TTM-Base ekambaram2024tiny as the base model. The training datasets includes 38.7 million time points from multiple public benchmarks. Evaluation is conducted on the held-out ETT and Weather datasets, ensuring strict zero-shot conditions without data leakage. Detailed information can be found in Figure 9 of Appendix.
Single-source transfer zero-shot setup (for LLM based TSPMs) This setup trains a model on a single source dataset (e.g., ETTh1) and evaluates it on an unseen target dataset (e.g., ETTm2), highlighting the model’s ability to transfer knowledge across domains. Commonly used in LLM based TSPMs jin2023time ; zhou2023one , this setup involves fine-tuning or prompting LLMs on a specific dataset and then applying them to novel inputs from a different domain. We use this setup for QuiZSFL, with TimeLLM jin2023time as the base model and LLaMA-7B touvron2023llama as the backbone. Prompts are constructed with MSIL and domain-specific patterns, and evaluation is conducted across multiple train–test splits according to the TimeLLM ZSF experimental setup (Table 2).
5.3 Peformance Comparison
The prediction results of QuiZSFT and QuiZSFL are shown in Table 1 and Table 2, respectively. The best performance is marked in bold, and the second best is underlined.
QuiZSFT. To provide a more comprehensive assessment, we compare QuiZSFT with several strong full-shot forecasting models. These comparisons fall under the category of long-sequence forecasting. As shown in Table 1, "zero-shot" refers to the prediction results of various base models without any pre-training on the test datasets, while "full-shot" denotes the performance of benchmark models that have been fully trained on each dataset. QuiZSFT, trained solely under the zero-shot setting, outperforms not only existing zero-shot baselines but also full-shot models, demonstrating strong generalization capabilities even without access to target-domain training data. QuiZSFT ranks Top1 in 75% of ZSF settings. However, we observe that QuiZSFT performs particularly well on relatively coarse-grained datasets but shows limited effectiveness on short-term, minute-scale forecasting tasks (e.g., ETTm1). This may be due to the difficulty in aligning retrieved information with fine-grained fluctuations in the target sequence. Leveraging coarse-grained knowledge to guide fine-grained prediction presents an interesting direction for future research. Combining complexity and efficiency comparison in Figure 4 (a), we consider QuiZSFT to be an excellent lightweight zero-shot forecasting framework, which benefits from the retrieval-enhanced representation and active feedback.
To further validate the generalization capability of QuiZSFT, we also evaluate its performance when the domain of the target series is not within the seven domains of CRB. For this purpose, we use the Weather dataset (Meteorological Domain) as the test set for experimentation, with results shown in Table 1. QuiZSFT achieves the best performance across all prediction horizons, demonstrating its strong generalization capability.
Table 1: Long sequence forecasting results. The experimental setup follows TTM ekambaram2024tiny . Best results are in bold; second best are underlined. Full-shot results are obtained from the Moirai woo2024unified where the authors draw similar comparison.
Zero-shot forecasting Full-shot forecasting
QuiZTST
TTMB
MoiraiB
TimesFM
iTransformer
Crossformer
DLinear
TimesNet
PatchTST
TiDE
FEDformer
ETTh1 96 0.361 0.364 0.384 0.421 0.386 0.423 0.386 0.384 0.414 0.479 0.376
192 0.384 0.388 0.425 0.472 0.441 0.471 0.437 0.436 0.46 0.525 0.420
336 0.398 0.402 0.456 0.51 0.487 0.570 0.481 0.491 0.501 0.565 0.459
720 0.468 0.471 0.470 0.514 0.503 0.653 0.519 0.521 0.500 0.594 0.506
Avg 0.403 0.406 0.434 0.479 0.454 0.529 0.456 0.458 0.469 0.541 0.440
ETTh2 96 0.276 0.279 0.277 0.326 0.297 0.745 0.333 0.34 0.302 0.4 0.358
192 0.334 0.334 0.34 0.400 0.380 0.877 0.477 0.402 0.388 0.528 0.429
336 0.364 0.366 0.371 0.434 0.428 1.043 0.594 0.452 0.426 0.643 0.496
720 0.407 0.408 0.394 0.451 0.427 1.104 0.831 0.462 0.431 0.874 0.463
Avg 0.345 0.347 0.346 0.403 0.383 0.942 0.559 0.414 0.388 0.611 0.437
ETTm1 96 0.369 0.359 0.335 0.357 0.334 0.404 0.345 0.338 0.329 0.364 0.379
192 0.377 0.376 0.367 0.411 0.337 0.450 0.38 0.374 0.367 0.398 0.426
336 0.397 0.407 0.398 0.442 0.426 0.532 0.413 0.41 0.409 0.428 0.445
720 0.441 0.446 0.434 0.507 0.491 0.666 0.474 0.478 0.481 0.487 0.543
Avg 0.395 0.397 0.383 0.429 0.397 0.513 0.403 0.400 0.397 0.419 0.448
ETTm2 96 0.176 0.178 0.195 0.205 0.18 0.287 0.193 0.187 0.175 0.207 0.203
192 0.238 0.238 0.247 0.293 0.25 0.414 0.284 0.249 0.241 0.290 0.269
336 0.292 0.300 0.293 0.366 0.311 0.597 0.369 0.321 0.305 0.377 0.325
720 0.390 0.41 0.365 0.472 0.412 1.730 0.554 0.408 0.402 0.558 0.421
Avg 0.274 0.282 0.275 0.334 0.288 0.757 0.350 0.291 0.281 0.358 0.305
Weather 96 0.153 0.158 0.167 - 0.174 0.158 0.196 0.172 0.177 0.202 0.217
192 0.194 0.206 0.209 - 0.221 0.206 0.237 0.219 0.225 0.242 0.276
336 0.251 0.260 0.256 - 0.278 0.272 0.283 0.280 0.278 0.287 0.339
720 0.324 0.330 0.325 - 0.358 0.398 0.345 0.365 0.354 0.351 0.403
Avg 0.231 0.239 0.239 - 0.258 0.259 0.265 0.259 0.259 0.271 0.309
Table 2: Zero-shot forecasting results under the single-source transfer setup. Following the setting of TimeLLM jin2023time , results are averaged over prediction lengths {96, 192, 336, 720}. Best scores are in bold, second best are underlined.
QuiZSFL Time-LLM LLMTime GPT4TS DLinear PatchTST TimesNet Autoformer
ETTh1 $\rightarrow$ ETTh2 0.352 0.356 0.992 0.406 0.493 0.380 0.421 0.582
ETTh1 $\rightarrow$ ETTm2 0.272 0.277 1.867 0.325 0.415 0.314 0.327 0.457
ETTh2 $\rightarrow$ ETTh1 0.535 0.521 1.961 0.757 0.703 0.545 0.865 0.757
ETTh2 $\rightarrow$ ETTm2 0.269 0.271 1.867 0.335 0.328 0.325 0.342 0.366
ETTm1 $\rightarrow$ ETTh2 0.382 0.394 0.992 0.433 0.464 0.439 0.457 0.470
ETTm1 $\rightarrow$ ETTm2 0.281 0.296 1.867 0.313 0.335 0.291 0.322 0.469
ETTm2 $\rightarrow$ ETTh2 0.351 0.354 0.992 0.435 0.455 0.409 0.435 0.423
ETTm2 $\rightarrow$ ETTm1 0.414 0.418 1.993 0.769 0.649 0.568 0.769 0.755
QuiZSFL. As shown in Table 2, QuiZSFL, equipped with retrieved augmented series, outperforms the majority of competing methods, achieving the best results in 7 out of 8 prediction settings. Additionally, our approach enhances the performance based on the pre-trained model. With the continuous advancement of pre-trained models, it is anticipated that QuiZSFL will bring about further performance improvements when adapting to new model frameworks.
5.4 Ablation study
Ablative variants. 1) QuiZSFL-w/o-RAG. We remove the auxiliary sequence retrieval and only utilize the LLM for ZSF to verify the motivation for RAG, which degenerates to LLMTime (gruver2024large, ). 2) QuiZSFL-w/o-MSIL. We remove the Multi-grained Series Interaction Learner, performing only an average calculation on the retrieved time series instead of feature extraction. 3) QuiZSFL-w/o-Coherer. We remove the structured prompt template and directly concatenate MSIL-extracted features with the target sequence, without adapting them into LLM-compatible textual inputs.
Main results. Results are in Table 3. The most significant performance drop occurs when removing the retrieval module (RAG), confirming its importance in providing external contextual signals for zero-shot forecasting. Compared to the full model, this variant sees a performance decline of 3.14%–5.34% (line 1 vs. line 4). Removing MSIL and using simple averaging instead also degrades performance. This is because the model cannot learn multi-level representations and capture relationships between sequences, resulting in performance degradation (line 2 vs. line 4). Additionally, without converting numerical time series into LLM-understandable tokens using the prompt template, the LLM’s performance is inferior to the full QuiZSFL, with a drop of about 2% (line 3 vs. line 4). These results demonstrate the effectiveness of our integrated approach.
Table 3: Ablation studies on zero-shot settings
ETTm1 → ETTh2 ETTm1 → ETTm2
QuiZSFL-w/o-RAG 0.394 0.296
QuiZSFL-w/o-MSIL 0.387 0.286
QuiZSFL-w/o-Coherer 0.388 0.289
QuiZSFL 0.382 0.281
5.5 Detailed model analysis
Complexity analysis. We report empirical comparisons of model size and inference time in Figure 4 (a). QuiZSFT maintains competitive efficiency while outperforming most baselines. Although it introduces minor computational overhead due to retrieval and interaction modules, the added cost is minimal. Full analysis is provided in Appendix D.1.
Hyperparameter analysis. Key hyperparameters include the size of the retrieval database, the number of retrieved sequences $K$, and the local-domain proportion $\rho$. As shown in Figure 4 (b), performance is sensitive to these choices. CRB-Medium, $K=8$, and $\rho=60\%$ strike a good balance between accuracy and efficiency. Detailed experiments and discussion are provided in Appendix D.2.
Figure 4: (a) Size and time overview of QuiZTST vs. pre-trained TS benchmarks. Plot each model based on model size and the CPU inference time per batch. (b) Hyperparameter analysis.
Case study. We visualize results on the ETTh2 dataset in Figure 5. The retrieved patterns help smooth out noisy predictions and reduce hallucinations. A complete analysis of sequence selection and LLM behavior is included in Appendix D.3.
Figure 5: Case studies on ETTh2 prediction.
6 Conclusion
We present QuiZSF, a lightweight and efficient framework that integrates TSPMs with RAG for zero-shot time series forecasting. By combining tree-structured base construction, hierarchical retrieval, multi-level representation learning, and modality-aware integration, QuiZSF effectively utilizes both inter- and intra-domain sequences to enhance forecasting accuracy. QuiZSF achieves strong performance in both zero-shot and full-shot settings and offers practical insights for time series modeling in data-scarce scenarios such as domain transfer and extreme conditions. It also holds promise for broader applications in intelligent data systems and temporal databases.
Appendix A Detailed Design of ChronoRAG Base
A.1 Data protocol for ChronoRAG
Large-scale time series datasets are essential for retrieval tasks, but constructing the ChronoRAG Base (CRB) presents challenges such as inconsistent lengths, dimensionality mismatches, missing values, metadata diversity, and storage scalability. To tackle these issues, we design a unified data protocol that ensures consistent preprocessing, metadata unification, and efficient storage. All data is stored using the ARROW format peltenburg2019fletcher , which is optimized for deep learning frameworks and enables efficient retrieval and access.
To address the issue of inconsistent data lengths across datasets, we adopt a sliding window approach. Let $\mathbf{X}\in\mathbb{R}^{N\times D}$ be a multivariate time series with $N$ time steps and $D$ channels. The $j$-th channel is denoted as $\mathbf{x}_{j}=[x_{1j},x_{2j},\cdots,x_{Nj}]$. Following the foundation model settings ekambaram2024tiny , we set window size $w$ and step size $s$ ($w\leq N$), and segment each channel into uniform windows:
$\mathbf{X}_{kj}=[x_{(k-1)s+1,j},\cdots,x_{(k-1)s+w,j}]$ (15)
where $(k-1)s+w\leq N$. This preserves local patterns and improves retrieval efficiency.
To address the issue of varying dimensionality, we use a channel-independent strategy, which treats each dimension separately and has been validated by PatchTST nie2022time , TimeLLM jin2023time , and TTM ekambaram2024tiny . For each channel $\mathbf{x}_{i}$, we apply a shared function $f$:
$\mathbf{y}_{i}=f(\mathbf{x}_{i})$ (16)
This simplifies database construction and fusion, while enhancing scalability and cross-domain adaptability.
To address the issue of missing values, which may impair data integrity and affect retrieval, we apply linear interpolation friedman1962interpolation to complete incomplete sequences.
To address the issue of diverse metadata, we define a unified metadata protocol by standardizing key attributes such as item ID, start time, end time, frequency, domain, and sequence values (see Table 4). This ensures consistent integration across multi-source datasets.
To address the issue of large volume and variety of sequence data, which make efficiently storing and retrieving a significant challenge, we implement a hierarchical tree-like storage structure berchtold1996x ; jensen2004query ; athanassoulis2014bf . This enables efficient storage and indexing for large-scale datasets and seamless integration into deep learning frameworks. Details are provided in Section 6.
Table 4: Structural key-value instance in CRB
Meta information Deterministic observation
Key
Domain
Category
Item_id Start End Freq Target
Value Nature
us_births_
dataset_0_0
20000101 20010527 Daily [9083,8006,11136,……]
A.2 Hierarchical Series Tree
Figure 6: Tree-shaped hierarchical series organization with pre-clustering. Prototypes represent cluster centroids.
To support efficient indexing and retrieval in ChronoRAG Base (CRB), we design a hierarchical tree structure with pre-clustering. Traditional FIFO-based linear storage pelkonen2015gorilla ; jensen2017time suffers from inefficiency when scaling to millions of time series. Linear retrieval requires one-by-one comparisons with time complexity $T_{\text{linear}}=O(N)$, which becomes prohibitive at large scale.
To mitigate this, we construct a tree-shaped structure inspired by database indexing techniques berchtold1996x ; jensen2004query ; athanassoulis2014bf . At the top level, the database is partitioned by domain: given a dataset $\mathcal{X}=\{X_{1},\dots,X_{N}\}$, we separate it into $K$ disjoint domain-based groups $\{\mathcal{D}_{k}\}_{k=1}^{K}$, such that:
$\mathcal{X}=\bigcup_{k=1}^{K}\mathcal{D}_{k},\quad\mathcal{D}_{i}\cap\mathcal{D}_{j}=\emptyset\text{ for }i\neq j.$ (17)
Each domain group $\mathcal{D}_{k}$ is then recursively divided using the $k$-means algorithm, where each cluster contains at most $N=256$ time series. This setting follows the clustering granularity used in prior work such as Marigold mortensen2023marigold , where small prototype groups (typically $N\leq 256$) are shown to improve retrieval quality and update flexibility. The number of clusters $M_{k}$ in each domain is thus determined by the size of $\mathcal{D}_{k}$, i.e., $M_{k}=\lceil\frac{|\mathcal{D}_{k}|}{N}\rceil$.
Each cluster $\mathcal{C}_{m}^{(k)}$ is formed by minimizing the standard $k$-means objective:
$\mathcal{L}_{k}=\sum_{X_{i}\in\mathcal{D}_{k}}\|X_{i}-C_{j(i)}^{(k)}\|^{2},$ (18)
where $C_{j(i)}^{(k)}$ is the centroid of the cluster to which $X_{i}$ belongs. The prototype for each cluster is selected as the sequence closest to the centroid:
$X_{\text{proto}}^{(k,m)}=\arg\min_{X_{i}\in\mathcal{C}_{m}^{(k)}}\|X_{i}-C_{m}^{(k)}\|^{2}.$ (19)
While tree-based structures have a theoretical average-case complexity of $O(\log_{b}N)$, this does not always hold in high-dimensional time series due to the curse of dimensionality beyer1999nearest . Hence, instead of relying solely on theoretical claims, we report empirical improvements in retrieval speed and memory in Section 5.5.
This hierarchical prototype-based structure supports domain-level filtering and accelerates approximate nearest-neighbor search. During retrieval, a query is first matched against cluster prototypes, and then only a small number of candidate clusters are examined in full. This greatly reduces computation compared to flat comparisons.
Finally, the tree structure supports dynamic updates. When a new sequence $X_{\text{new}}$ arrives, it is first matched to the nearest prototype using:
$X_{\text{proto}}^{*}=\arg\min_{X_{\text{proto}}}\|X_{\text{new}}-X_{\text{proto}}\|^{2},$ (20)
and inserted into the corresponding cluster $\mathcal{C}_{m}^{(k)}$. If the cluster exceeds the predefined maximum size $N$, local re-clustering is triggered within the affected subtree.
This local re-clustering process reassigns the sequences in the overflowed cluster by minimizing the intra-cluster distance:
$\min_{\{\mathcal{C}_{i}^{(k)}\}}\sum_{X_{j}\in\bigcup\mathcal{C}_{i}^{(k)}}\left\|X_{j}-C_{i(j)}^{(k)}\right\|^{2},$ (21)
where the optimization is restricted to the current subtree, and $C_{i(j)}^{(k)}$ denotes the centroid of the sub-cluster assigned to $X_{j}$. This ensures that updates remain computationally tractable and localized.
The above update strategy allows the index to evolve incrementally over time without full reorganization, and is inspired by dynamic clustering methods in data streams cao2006density ; aghabozorgi2015time .
A.3 Database Versions
The three versions of CRB are illustrated in Figure 7. Each version covers all seven domains, and the smaller versions are subsets of the larger ones. Detailed information about CRB-Large can be found in Table 5.
Figure 7: Three versions of CRB.
Table 5: CRB-Large Detailed Descriptions: Domain indicates the field to which the dataset belongs. Datasets refer to the specific datasets included. Time Series represents the number of time series contained in the dataset after processing. Frequency denotes the sampling interval of time points, where “-” indicates either the absence of timestamps or irregular intervals. Time Points represents the total number of time points in the dataset. Source specifies the original paper or resource from which the dataset is obtained.
Domain Datasets Time Series Frequency Time Points Source
Energy wind_4_seconds_dataset 1 4 Sec 512 Monash godahewa2021monash
australian_electricity_demand_dataset 5 30 Min 2560 Monash godahewa2021monash
london_smart_meters_dataset_without_missing_values 5556 Hourly 2844672 Monash godahewa2021monash
Health SelfRegulationSCP1 3366 0.004 Sec 1723392 UCR Time Series Archive dau2019ucr
MotorImagery 24192 0.001 Sec 12386304 UCR Time Series Archive dau2019ucr
PigCVP 312 - 159744 UCR Time Series Archive dau2019ucr
PigArtPressure 312 - 159744 UCR Time Series Archive dau2019ucr
SelfRegulationSCP2 2660 0.004 Sec 1361920 UCR Time Series Archive dau2019ucr
AtrialFibrillation 60 0.008 Sec 30720 UCR Time Series Archive dau2019ucr
IEEEPPG 15480 0.008 Sec 7925760 TSER archive tan2021time
BIDMC32HR 12278 - 6286336 TSER archive tan2021time
TDBrain 28644 0.002 Sec 14665728 TDBrain van2022two
IoT baian 918 0.02 Sec 470016 UTSD  liutimer
Nature StarLightCurves 9236 - 4728832 UCR Time Series Archive dau2019ucr
Phoneme 2110 - 1080320 UCR Time Series Archive dau2019ucr
EigenWorms 1554 - 795648 UCR Time Series Archive dau2019ucr
Worms 258 0.033 Sec 132096 UCR Time Series Archive dau2019ucr
us_births_dataset 1 Daily 512 Monash godahewa2021monash
kdd_cup_2018_dataset_without_missing_values 270 Hourly 138240 Monash godahewa2021monash
temperature_rain_dataset_without_missing_values 32072 Daily 16420864 Monash godahewa2021monash
Sunspot_dataset_without_missing_values 1 Daily 512 Monash godahewa2021monash
saugeenday_dataset 1 Daily 512 Monash godahewa2021monash
Web kaggle_web_traffic_dataset_without_missing_values 141444 Daily 72419328 Monash godahewa2021monash
Transport pedestrian_counts_dataset 66 Hourly 33792 Monash godahewa2021monash
Environment AustraliaRainfall 3 Hourly 1536 TSER archive tan2021time
BenzeneConcentration 8 Hourly 4096 TSER archive tan2021time
BeijingPM25Quality 9 Hourly 4608 TSER archive tan2021time
Appendix B Language Coherer for LLM based TSPMs
Unlike numerical models, language models operate exclusively on text. Direct integration of numeric features is thus infeasible due to modality mismatch. To bridge this gap, we transform the MSIL-derived representations (i.e., $\mathbf{P}_{\text{int}}$, $\mathbf{P}_{\text{avg}}$, and $\mathbf{T}^{\text{norm}}$) into structured textual summaries. These summaries are combined with an instruction prompt to guide the language models in generating forecasting outputs.
The transformation process involves converting numerical features into a format that can be understood by language models. This is achieved by creating textual summaries that capture the essential characteristics of the retrieved time series. The summaries are then combined with an instruction prompt that provides context and guidance for the language model to generate accurate forecasts.
This approach ensures that the language model can effectively leverage the retrieved knowledge, even though it operates on a different modality. By converting numerical features into structured text, we enable seamless integration with the language model, allowing it to generate more accurate and reliable forecasts.
Figure 8: Prompt construction framework for LLM-based forecasting.
Appendix C Training Dataset Details
Detailed information of the training set is shown in Figure 9. The x-axis denotes domains. In addition to the seven core domains (Table 5), extra domains are included to enhance generalization. The y-axis shows the number of time points and datasets per domain.
Figure 9: Detailed information of the training set.
Appendix D Additional Experimental Analyses
D.1 Complexity Analysis
We conduct a detailed empirical analysis and present a comparison of the complexity and efficiency of the QuiZSFT in Figure 4 (a). The comparison is mainly carried out from two aspects: the model size (in mebibytes, MiB) and the CPU inference time per batch (s/iter). QuiZSFT shows a notable performance in both metrics, with a clear advantage over most comparative models, falling just slightly short of TTMB. Upon further investigation, it is found that this is due to the introduction of learnable modules in the retrieval and feature extraction process based on Retrieval-Augmented Generation (RAG). However, it is worth mentioning that these modules introduced by QuiZSFT are of a lightweight design, with very few additional parameters. Moreover, the retrieval and feature extraction processes rely on dot product calculations, which are highly efficient and do not significantly extend inference time. The empirical results clearly indicate that while the introduction of QuiZSFT brings certain memory and time overheads, these overheads are within an acceptable range, and at the same time, the model’s performance is significantly enhanced. We further believe that by properly adjusting the hyperparameters in the retrieval process and the learnable weights, the number of model parameters can be reduced, thus further optimizing the model.
D.2 Hyperparameter Study and Analysis
We specify three key hyperparameters to explore how to achieve their optimal performance on QuiZSFL. Firstly, "Database Scale", which includes three scales of CRB, namely {Small, Medium, Large}. Secondly, it is the number of time series retrieved from the auxiliary sequence dataset, with values in the range of $K=\{1,4,8,12,16\}$. Thirdly, it is the proportion of retrieved time series that are in the same domain as the target series, that is, the Local prototype $\rho$, with values of $\{0\%,20\%,40\%,60\%,80\%,100\%\}$. Due to space limitations, the first two experiments are only elaborated in the ETTm1 $\rightarrow$ ETTm2 task, while the last experiment is described in both the ETTm1 $\rightarrow$ ETTm2 and ETTm1 $\rightarrow$ ETTh2 tasks, as shown in Figure 10.
The CRB_Small scale performs worst (Figure 10 (a)). As the scale decreases, the external knowledge it provides declines, leading to poor performance. This partly verifies the scaling law kaplan2020scaling in time series. In the hyperparameter experiment for the retrieved number K, TopK =8=8 yields the best results. Retrieving more sequences may introduce more noise, while fewer sequences carry less information (Figure 10 (b)). In the hyperparameter experiment for the Local prototype $\rho$, two tasks show different trends. A too-high proportion of the same domain limits data feature diversity, over-emphasizing single-dimension features. The model performs well on samples fitting this feature (Figure 10 (c)), but poorly on non-matching samples due to the lack of auxiliary correction from other dimensions, resulting in a polarized outcome (Figure 10 (d)). To balance performance and efficiency, we select CRB_Medium, set TopKK to 8, and set the Local prototype $\rho$ at $60\%$.
Figure 10: Hyperparameter analysis.
D.3 Case Study
In order to demonstrate how the retrieval sequence improves the prediction effect, we conduct an intuitive analysis of the intermediate results. Taking the target sequence of the ETTh2 dataset as an example (as shown in Figure 5), we screen out the Top-8 sequences that are closest to the target sequence through hybrid and hierarchical time-series retrieval and marked their similarities. These sequences have similar patterns and evolution models as the target sequence. Subsequently, we update the sequence representation of the target sequence by combining the auxiliary sequence with the target sequence through MSIL and input it into the LLM after generating prompts. We visualized and compared the output of the LLM with RAG and the output w/o RAG.
The results show that RAG can reveal the average pattern of the retrieval sequence, making the prediction results smoother and avoiding overfitting; while the output w/o RAG fluctuates more and contains more inaccurate details. This indicates that RAG effectively suppresses the time-series hallucination of the LLM. Our analysis enhances the interpretability of the model, deepens the understanding of zero-shot forecasting, and highlights the contribution of RAG in enhancing prediction.
D.4 Experiments compute resources
All experiments were conducted on a single NVIDIA A100 GPU with 40GB memory.
Appendix E Limitations and Future Work
While our proposed framework, QuiZSF, demonstrates significant advancements in zero-shot forecasting for time series data, there are several limitations that warrant further investigation. First, the effectiveness of QuiZSF is tied to the quality and relevance of pre-trained models, which can be a limiting factor in domains with limited or highly specialized data. Future work could explore the development of more robust and domain-specific pre-trained models to enhance the framework’s capabilities. Second, reducing the complexity of the framework is crucial for large-scale retrieval processes. Introducing sparse connections into the retrieval database and exploring such connections to facilitate relevance retrieval could significantly improve efficiency. Lastly, developing multi-granularity sequential learning methods to allow pattern transfer between different sub-sequences could address cross-granularity transfer challenges. These advancements would further enhance the scalability and applicability of QuiZSF in diverse forecasting scenarios.